# Modeling Item-Level Dynamic Variability with Residual Diffusion for Bundle Recommendation

### v2

@@ 3.2. Overall Performance (RQ1)
$\rho$ 4 2 0 -2 -4 4 2 0 -2 -4
Methods R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20
BGCN 0.1897 0.1047 0.2254 0.1278 0.2367 0.1380 0.2214 0.1247 0.2138 0.1204 0.0581 0.0309 0.0634 0.0333 0.0643 0.0339 0.0646 0.0341 0.0622 0.0332
+RDiffBR 0.2178 0.1278 0.2391 0.1415 0.2453 0.1466 0.2470 0.1423 0.2407 0.1396 0.0614 0.0321 0.0657 0.0340 0.0662 0.0343 0.0668 0.0345 0.0649 0.0340
improv. 14.83% 22.15% 6.06% 10.74% 3.61% 6.25% 11.57% 14.11% 12.60% 16.03% 5.56% 3.98% 3.63% 1.99% 2.89% 1.14% 3.47% 1.07% 4.42% 2.54%
MIDGN 0.2388 0.1373 0.2557 0.1472 0.2598 0.1496 0.2578 0.1495 0.2566 0.1489 0.0540 0.0278 0.0625 0.0330 0.0636 0.0337 0.0639 0.0337 0.0640 0.0337
+RDiffBR 0.2498 0.1441 0.2621 0.1498 0.2645 0.1515 0.2652 0.1517 0.2643 0.1516 0.0568 0.0293 0.0649 0.0340 0.0661 0.0347 0.0667 0.0348 0.0672 0.0350
improv. 4.62% 4.98% 2.53% 1.72% 1.79% 1.32% 2.87% 1.49% 3.00% 1.82% 5.21% 5.27% 3.89% 2.91% 3.96% 2.92% 4.45% 3.25% 4.95% 4.14%
HyperMBR 0.1464 0.0953 0.2126 0.1290 0.2649 0.1581 0.2167 0.1285 0.1897 0.1162 0.0588 0.0313 0.0675 0.0354 0.0717 0.0377 0.0689 0.0363 0.0661 0.0345
+RDiffBR 0.1789 0.1109 0.2229 0.1327 0.2669 0.1594 0.2280 0.1347 0.2128 0.1275 0.0627 0.0332 0.0694 0.0369 0.0739 0.0386 0.0704 0.0369 0.0667 0.0350
improv. 22.20% 16.42% 4.85% 2.87% 0.75% 0.83% 5.20% 4.87% 12.14% 9.75% 6.67% 5.97% 2.88% 4.41% 3.07% 2.51% 2.25% 1.79% 0.85% 1.50%
CrossCBR 0.2677 0.1584 0.2763 0.1626 0.2776 0.1640 0.2788 0.1649 0.2800 0.1656 0.0715 0.0386 0.0776 0.0419 0.0794 0.0431 0.0802 0.0436 0.0798 0.0434
+RDiffBR 0.2744 0.1640 0.2814 0.1659 0.2827 0.1663 0.2816 0.1664 0.2816 0.1662 0.0772 0.0419 0.0818 0.0436 0.0826 0.0443 0.0829 0.0444 0.0826 0.0445
improv. 2.52% 3.57% 1.84% 2.02% 1.85% 1.42% 1.01% 0.92% 0.56% 0.40% 7.99% 8.50% 5.36% 4.04% 4.09% 2.67% 3.43% 1.70% 3.50% 2.51%
MultiCBR 0.2770 0.1649 0.2769 0.1649 0.2791 0.1652 0.2767 0.1641 0.2770 0.1648 0.0843 0.0463 0.0867 0.0472 0.0876 0.0475 0.0880 0.0479 0.0879 0.0479
+RDiffBR 0.2831 0.1674 0.2840 0.1677 0.2824 0.1673 0.2825 0.1669 0.2842 0.1675 0.0859 0.0469 0.0879 0.0478 0.0881 0.0480 0.0883 0.0482 0.0893 0.0487
improv. 2.20% 1.55% 2.56% 1.71% 1.19% 1.25% 2.10% 1.71% 2.60% 1.61% 1.91% 1.32% 1.37% 1.27% 0.57% 1.18% 0.40% 0.75% 1.60% 1.52%
BundleGT 0.2687 0.1590 0.2872 0.1683 0.2877 0.1694 0.2873 0.1697 0.2837 0.1679 0.0729 0.0367 0.0869 0.0465 0.0874 0.0468 0.0874 0.0468 0.0871 0.0465
+RDiffBR 0.2829 0.1653 0.2927 0.1708 0.2922 0.1707 0.2939 0.1712 0.2903 0.1702 0.0759 0.0384 0.0900 0.0479 0.0896 0.0479 0.0897 0.0477 0.0895 0.0476
improv. 5.27% 3.96% 1.90% 1.49% 1.57% 0.79% 2.30% 0.93% 2.31% 1.40% 4.06% 4.75% 3.49% 3.16% 2.52% 2.42% 2.67% 1.86% 2.76% 2.37%
Dataset iFashion $\mathbf{{MealRec+}}$
$\rho$ 4 2 0 -2 -4 4 2 0 -2 -4
Methods R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20 R@20 N@20
BGCN 0.0473 0.0331 0.0560 0.0395 0.0602 0.0426 0.0607 0.0428 0.0604 0.0427 0.1068 0.1888 0.1213 0.2094 0.1325 0.2247 0.1326 0.2250 0.1375 0.2323
+RDiffBR 0.0488 0.0341 0.0568 0.0398 0.0606 0.0428 0.0608 0.0429 0.0607 0.0429 0.1180 0.2078 0.1293 0.2180 0.1408 0.2401 0.1387 0.2371 0.1388 0.2384
improv. 3.09% 2.91% 1.55% 0.58% 0.61% 0.45% 0.16% 0.23% 0.57% 0.47% 10.07% 10.55% 4.08% 6.53% 6.82% 6.26% 5.38% 4.57% 2.63% 0.92%
MIDGN 0.0372 0.0257 0.0549 0.0380 0.0672 0.0482 0.0428 0.0279 0.0441 0.0294 0.0820 0.1564 0.1017 0.1816 0.1121 0.1908 0.1129 0.1919 0.1115 0.1880
+RDiffBR 0.0426 0.0290 0.0628 0.0437 0.0763 0.0553 0.0534 0.0348 0.0545 0.0363 0.0826 0.1604 0.1075 0.1930 0.1179 0.2050 0.1145 0.2012 0.1146 0.2027
improv. 14.55% 12.83% 14.36% 14.87% 13.53% 14.65% 24.66% 24.83% 23.48% 23.77% 2.60% 0.82% 6.29% 5.69% 7.44% 5.13% 4.87% 1.48% 7.82% 2.79%
HyperMBR 0.0448 0.0336 0.0687 0.0494 0.0815 0.0580 0.0773 0.0545 0.0744 0.0521 0.0859 0.1498 0.0939 0.1595 0.1069 0.1778 0.1057 0.1781 0.1106 0.1887
+RDiffBR 0.0462 0.0349 0.0692 0.0504 0.0830 0.0595 0.0797 0.0567 0.0773 0.0550 0.0874 0.1585 0.0995 0.1760 0.1122 0.1944 0.1205 0.1988 0.1247 0.2144
improv. 3.21% 4.09% 0.81% 2.01% 1.75% 2.69% 3.01% 4.01% 3.92% 5.42% 5.82% 1.70% 10.33% 5.90% 9.34% 4.94% 11.61% 14.00% 13.62% 12.82%
CrossCBR 0.0624 0.0491 0.0877 0.0683 0.0999 0.0772 0.1046 0.0799 0.1081 0.0831 0.0612 0.1048 0.0819 0.1417 0.0938 0.1688 0.1084 0.2041 0.1504 0.2599
+RDiffBR 0.0720 0.0554 0.0928 0.0716 0.1027 0.0794 0.1060 0.0812 0.1088 0.0834 0.0681 0.1203 0.0926 0.1531 0.1026 0.1788 0.1241 0.2243 0.1611 0.2669
improv. 15.36% 12.87% 5.75% 4.78% 2.83% 2.82% 1.31% 1.64% 0.60% 0.32% 14.80% 11.40% 8.01% 13.15% 5.95% 9.37% 9.89% 14.48% 2.69% 7.12%
MultiCBR 0.0672 0.0588 0.1080 0.0905 0.1297 0.1064 0.1343 0.1101 0.1381 0.1133 0.0989 0.1782 0.1147 0.1841 0.1305 0.2092 0.1476 0.2376 0.1625 0.2592
+RDiffBR 0.0694 0.0608 0.1094 0.0916 0.1313 0.1080 0.1350 0.1111 0.1385 0.1139 0.1066 0.1889 0.1217 0.1945 0.1375 0.2164 0.1524 0.2436 0.1633 0.2598
improv. 3.29% 3.31% 1.23% 1.16% 1.20% 1.47% 0.53% 0.97% 0.27% 0.58% 6.04% 7.83% 5.63% 6.10% 3.44% 5.40% 2.51% 3.19% 0.25% 0.50%
BundleGT 0.0444 0.0352 0.0690 0.0517 0.0954 0.0724 0.0899 0.0653 0.0960 0.0707 0.0732 0.1281 0.0923 0.1556 0.1093 0.1761 0.1217 0.2134 0.1579 0.2585
+RDiffBR 0.0456 0.0361 0.0701 0.0530 0.0970 0.0740 0.0909 0.0660 0.0970 0.0712 0.0863 0.1512 0.1082 0.1770 0.1315 0.2077 0.1414 0.2346 0.1707 0.2733
improv. 2.68% 2.58% 1.64% 2.34% 1.61% 2.10% 1.02% 1.04% 1.10% 0.71% 18.03% 17.88% 13.75% 17.20% 17.97% 20.35% 9.91% 16.21% 5.73% 8.05%
We present the results of Recall@20 and NDCG@20 for five levels of B-I variations in Table 2. We show the results of Recall@40 and NDCG@40 on two datasets for ten levels of B-I variations in Figure 4. The following observations can be made:
@@
$\bullet$ The perfomance of BR models when meeting item-level dynamic variability. When $\rho<0$ ($\mathbf{Z_{test}}$ removes $\Delta^{-}$), the performance of bundle models is generally lower than that under $\rho=0$ in most cases. The greater the reduction in B-I, the more significant the performance degradation. When $\rho>0$ ($\mathbf{Z_{test}}$ adds $\Delta^{+}$), some models exhibit performance decline, others maintain comparable performance, while a few show improvement. This indicates that BR models possess capability to handle B-I variations to some extent, yet they tend to fail under drastic reductions of B-I. Notably, HyperMBR performs the worst, where both B-I adding and removing frequently lead to performance deterioration. This phenomenon may stem from hyperbolic space embeddings being more sensitive to distance errors compared to Euclidean space. In contrast, models like MIDGN, CrossCBR, and MultiCBR exhibit relatively stable performance, potentially attributed to their adoption of contrastive learning that leverages multi-view information through cooperative learning. Additionally, BundleGT demonstrates enhanced adaptability for dynamic B-I variations by modeling B-I affiliation via attention mechanisms.
@@
Figure 4. Performance comparisons of the backbone models and our RDiffBR under different levels of B-I variations in terms of Recall@40 and NDCG@40.
3.3. Ablation Study (RQ2)
@@
Table 3. Ablation study where bold indicates the best method, and underlined indicates the second best method($\rho=-3,3$).
Method Youshu NetEase $\mathrm{MealRec+}$
@@
Figure 5. Blue is distribution $E^{IL}_{B}$ under $\rho=0$. Red is distribution $E^{IL}_{B}$ under $\rho=-3$.
Figure 6. Blue is distribution $E^{IL}_{B}$ under $\rho=0$. Red is distribution $E^{IL}_{B}$ under $\rho=3$.
@@
From this table, we can draw the following conclusions: In most cases, the performance of using RDiffBR is superior to that of not using it or using RDiffBR w/o R. Using RDiffBR w/o R may lead to worse performance than RDiffBR, and in some cases it may even be worse than only using backbone BR model. Figure 6-6 show the visualization results of CrossCBR, CrossCBR+RDiffBR and CrossCBR+RDiffBR w/o R under two B-I varation levels for item-level bundle embeddings $E^{IL}_{B}$ in Youshu dataset. It can be observed that $E^{IL}_{B}$ generated by CrossCBR after B-I change is more dispersed than original $E^{IL}_{B}$ and more compact $E^{IL}_{B}$ distribution which retains the basic shape can be obtained after using CrossCBR+RDiffBR. This indicates that $E^{IL}_{B}$ of the BR model cannot effectively represent the bundle theme distribution due to the B-I change and RDiffBR enables the model to obtain more bundle information at the item-level rather than overfitting to the existing b-I affiliations. While using CrossCBR+RDiffBR w/o R, $E^{IL}_{B}$ distribution changes little so the effect is limited, which is consistent with the experimental results obtained in Table 3. We also conducted a case study to explore the specific situations of RDiffBR and RDiffBR w/o R(see detail in section3.5). Overall, RDiffBR outperforms backbone models, highlighting the effectiveness of RDiffBR. Residual connection plays an important role in RDiffBR, and merely using a diffusion model cannot generate better item-level bundle embeddings.
@@
Furthermore, our task can be approached without considering the Out-Of-Distribution(OOD) which means the distribution of the training data and the test data is inconsistent. Although different B-I affiliations are used during model testing, the bundle theme remains, which means items within a bundle remain to follow the similar distribution before and after the change. This can also be verified by Figure 6-6 that the shapes of two $E^{IL}_{B}$ distributions under different B-I variation $\rho$ are similar.
3.4. Impact of Hyperparameters
Figure 7. The impact of loss balance factor $\lambda$
Figure 8. The impact of residual weight $\delta$
Figure 9. The impact of dimensionality of time step embedding $d$
@@ 3.4. Impact of Hyperparameters
In order to achieve a more detailed understanding, we selected some hyperparameters: The loss balance factor $\lambda$, the residual weight $\delta$ and the dimensionality of time step embedding $d$. The results of these experiments are presented in Figure 9-9 by adjusting them within a certain range.
@@
$\bullet$ Loss balance factor $\lambda$. The loss balance factor $r$ affects the performance of RDiffBR, which is achieved by adjusting the emphasis between the recommendation task and the reconstruction task. A high value of $r$ may neglect the core recommendation task and thereby compromise the overall performance.
@@
$\bullet$ Residual weight $\delta$. The appropriate selection of residual weights is crucial for achieving embedding generation, this is because too high or too low values will lead to a decline in performance.
$\bullet$ Dimensionality of time step embedding $d$. As the value of d increases, the model performance initially improves. This is because the time step embedding can flexibly adjust the denoising strategy according to different noise stages. However, when the value of d is excessively large, the input embedding will gradually be weakened during the calculation process, thereby negatively affecting the model performance.
3.5. Case Study
@@
Figure 10. A case study for Youshu under $\rho=-3$. The highlighted items are those before the bundle change.
@@
Figure 11. A case study for Youshu under $\rho=3$. The highlighted items are those before the bundle change.
@@
In this section, we analyze the real data instances as show in Figure 11-11. We randomly select two bundles from Youshu and calculate cosine similarity between item-level bundle embeddings $E^{IL}_{B}$ and all items after B-I change. We listed the top 5 items with the highest similarity scores and their corresponding similarity values. It can be seen that the similarity of items(marked in red) that appeared within the bundle will be higher when only using CrossCBR, resulting in insufficient generalization ability. When using RDiffBR w/o R, the model cannot well capture the main items of the bundle. When using RDiffBR, the model can not only express the main items of the bundle but also introduce other items based on the theme to enhance the generalization ability.
3.6. Training Efficiency Study (RQ3)
@@
In order to evaluate the additional computing time of RDiffBR, we compared the one-epoch training time(seconds) of using BR model and BR model+RDiffBR on two different scale datasets. We record five consecutive training epochs and average them to obtain the one-epoch training time, as shown in Table 4. It shows that RDiffBR only increased training time about 4%.

### v3

@@ 13     Take gradient descent step on $\nabla_{\theta}(\mathcal{L})$ to optimize $\theta$;
$\rho$ -4 -2 0 2 4 -4 -2 0 2 4
Metrics R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20
BGCN .1897 .1047 .2254 .1278 .2367 .1380 .2214 .1247 .2138 .1204 .0581 .0309 .0634 .0333 .0643 .0339 .0646 .0341 .0622 .0332
+RDiffBR .2178 .1278 .2391 .1415 .2453 .1466 .2470 .1423 .2407 .1396 .0614 .0321 .0657 .0340 .0662 .0343 .0668 .0345 .0649 .0340
%improv. 14.83 22.15 6.06 10.74 3.61 6.25 11.57 14.11 12.60 16.03 5.56 3.98 3.63 1.99 2.89 1.14 3.47 1.07 4.42 2.54
MIDGN .2388 .1373 .2557 .1472 .2598 .1496 .2578 .1495 .2566 .1489 .0540 .0278 .0625 .0330 .0636 .0337 .0639 .0337 .0640 .0337
+RDiffBR .2498 .1441 .2621 .1498 .2645 .1515 .2652 .1517 .2643 .1516 .0568 .0293 .0649 .0340 .0661 .0347 .0667 .0348 .0672 .0350
%improv. 4.62 4.98 2.53 1.72 1.79 1.32 2.87 1.49 3.00 1.82 5.21 5.27 3.89 2.91 3.96 2.92 4.45 3.25 4.95 4.14
HyperMBR .1464 .0953 .2126 .1290 .2649 .1581 .2167 .1285 .1897 .1162 .0588 .0313 .0675 .0354 .0717 .0377 .0689 .0363 .0661 .0345
+RDiffBR .1789 .1109 .2229 .1327 .2669 .1594 .2280 .1347 .2128 .1275 .0627 .0332 .0694 .0369 .0739 .0386 .0704 .0369 .0667 .0350
%improv. 22.20 16.42 4.85 2.87 0.75 0.83 5.20 4.87 12.14 9.75 6.67 5.97 2.88 4.41 3.07 2.51 2.25 1.79 0.85 1.50
CrossCBR .2677 .1584 .2763 .1626 .2776 .1640 .2788 .1649 .2800 .1656 .0715 .0386 .0776 .0419 .0794 .0431 .0802 .0436 .0798 .0434
+RDiffBR .2744 .1640 .2814 .1659 .2827 .1663 .2816 .1664 .2816 .1662 .0772 .0419 .0818 .0436 .0826 .0443 .0829 .0444 .0826 .0445
%improv. 2.52 3.57 1.84 2.02 1.85 1.42 1.01 0.92 0.56 0.40 7.99 8.50 5.36 4.04 4.09 2.67 3.43 1.70 3.50 2.51
MultiCBR .2770 .1649 .2769 .1649 .2791 .1652 .2767 .1641 .2770 .1648 .0843 .0463 .0867 .0472 .0876 .0475 .0880 .0479 .0879 .0479
+RDiffBR .2831 .1674 .2840 .1677 .2824 .1673 .2825 .1669 .2842 .1675 .0859 .0469 .0879 .0478 .0881 .0480 .0883 .0482 .0893 .0487
%improv. 2.20 1.55 2.56 1.71 1.19 1.25 2.10 1.71 2.60 1.61 1.91 1.32 1.37 1.27 0.57 1.18 0.40 0.75 1.60 1.52
BundleGT .2687 .1590 .2872 .1683 .2877 .1694 .2873 .1697 .2837 .1679 .0729 .0367 .0869 .0465 .0874 .0468 .0874 .0468 .0871 .0465
+RDiffBR .2829 .1653 .2927 .1708 .2922 .1707 .2939 .1712 .2903 .1702 .0759 .0384 .0900 .0479 .0896 .0479 .0897 .0477 .0895 .0476
%improv. 5.27 3.96 1.90 1.49 1.57 0.79 2.30 0.93 2.31 1.40 4.06 4.75 3.49 3.16 2.52 2.42 2.67 1.86 2.76 2.37
[…]
$\rho$ -4 -2 0 2 4 -4 -2 0 2 4
Methods R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20 R20 N20
BGCN .0473 .0331 .0560 .0395 .0602 .0426 .0607 .0428 .0604 .0427 .1068 .1888 .1213 .2094 .1325 .2247 .1326 .2250 .1375 .2323
+RDiffBR .0488 .0341 .0568 .0398 .0606 .0428 .0608 .0429 .0607 .0429 .1180 .2078 .1293 .2180 .1408 .2401 .1387 .2371 .1388 .2384
%improv. 3.09 2.91 1.55 0.58 0.61 0.45 0.16 0.23 0.57 0.47 10.07 10.55 4.08 6.53 6.82 6.26 5.38 4.57 2.63 0.92
MIDGN .0372 .0257 .0549 .0380 .0672 .0482 .0428 .0279 .0441 .0294 .0820 .1564 .1017 .1816 .1121 .1908 .1129 .1919 .1115 .1880
+RDiffBR .0426 .0290 .0628 .0437 .0763 .0553 .0534 .0348 .0545 .0363 .0826 .1604 .1075 .1930 .1179 .2050 .1145 .2012 .1146 .2027
%improv. 14.55 12.83 14.36 14.87 13.53 14.65 24.66 24.83 23.48 23.77 2.60 0.82 6.29 5.69 7.44 5.13 4.87 1.48 7.82 2.79
HyperMBR .0448 .0336 .0687 .0494 .0815 .0580 .0773 .0545 .0744 .0521 .0859 .1498 .0939 .1595 .1069 .1778 .1057 .1781 .1106 .1887
+RDiffBR .0462 .0349 .0692 .0504 .0830 .0595 .0797 .0567 .0773 .0550 .0874 .1585 .0995 .1760 .1122 .1944 .1205 .1988 .1247 .2144
%improv. 3.21 4.09 0.81 2.01 1.75 2.69 3.01 4.01 3.92 5.42 5.82 1.70 10.33 5.90 9.34 4.94 11.61 14.00 13.62 12.82
CrossCBR .0624 .0491 .0877 .0683 .0999 .0772 .1046 .0799 .1081 .0831 .0612 .1048 .0819 .1417 .0938 .1688 .1084 .2041 .1504 .2599
+RDiffBR .0720 .0554 .0928 .0716 .1027 .0794 .1060 .0812 .1088 .0834 .0681 .1203 .0926 .1531 .1026 .1788 .1241 .2243 .1611 .2669
%improv. 15.36 12.87 5.75 4.78 2.83 2.82 1.31 1.64 0.60 0.32 14.80 11.40 8.01 13.15 5.95 9.37 9.89 14.48 2.69 7.12
MultiCBR .0672 .0588 .1080 .0905 .1297 .1064 .1343 .1101 .1381 .1133 .0989 .1782 .1147 .1841 .1305 .2092 .1476 .2376 .1625 .2592
+RDiffBR .0694 .0608 .1094 .0916 .1313 .1080 .1350 .1111 .1385 .1139 .1066 .1889 .1217 .1945 .1375 .2164 .1524 .2436 .1633 .2598
%improv. 3.29 3.31 1.23 1.16 1.20 1.47 0.53 0.97 0.27 0.58 6.04 7.83 5.63 6.10 3.44 5.40 2.51 3.19 0.25 0.50
BundleGT .0444 .0352 .0690 .0517 .0954 .0724 .0899 .0653 .0960 .0707 .0732 .1281 .0923 .1556 .1093 .1761 .1217 .2134 .1579 .2585
+RDiffBR .0456 .0361 .0701 .0530 .0970 .0740 .0909 .0660 .0970 .0712 .0863 .1512 .1082 .1770 .1315 .2077 .1414 .2346 .1707 .2733
%improv. 2.68 2.58 1.64 2.34 1.61 2.10 1.02 1.04 1.10 0.71 18.03 17.88 13.75 17.20 17.97 20.35 9.91 16.21 5.73 8.05
[…]
$\bullet$ BR Backbones. We implement RDiffBR on six representative BR models with official codes released by their authors. BGCN (BGCN) utilizes graph convolutional networks to separately learn the representations of bundle view and item view. HyperMBR (HyperMBR) models two views in hyperbolic space to learn their accurate representations. MIDGN (MIDGN) refines the user-item/bundle-item graph of different intents from multiple views. CrossCBR (CrossCBR) improves two separate views by constrastive learning. MultiCBR (MultiCBR) first fuses the multi-view representations into a unified one, and then adopts a contrastive loss on the unified representations. BundleGT (BundleGT) learns the strategy-aware user and bundle representation in bundle recommendation.
Figure 3: The processing of the B-I affiliation to match the scenario of B-I dynamic variability.
$\bullet$ Experimental Preprocess and Metrics. According to previous works (CrossCBR) (mealrec+), the training, validation and testing sets of user-bundle interactions $X$ are randomly splited with specific ratio. It is not feasible to simply add new bundles in bundle-item affiliations $Z$ under the B-I variability assumptions. Therefore, we adjust existing bundle datasets to match this scenario: Let $\{{\mathcal{E}_{1},\mathcal{E}_{2},…,\mathcal{E}_{10}}\}$ denote a uniform partition of $Z$ into ten disjoint subsets. Bundle models are trained and validated on $Z_{\mathrm{train\&val}}=\cup_{i=1}^{5}\mathcal{E}_{i}$. In testing phase, we design $Z_{\mathrm{test}}$ by adjusting $Z_{\mathrm{train\&val}}$ with adding or removing subsets:
$Z_{\mathrm{test}}=\begin{cases}Z_{\mathrm{train\&val}}\cup\left(\mathcal{\cup}_{i=6}^{5+\rho}\mathcal{E}_{i}\right)&\mathrm{if}~\rho>0\\ Z_{\mathrm{{train\&val}}}&\mathrm{if~\rho=0}\\ Z_{\mathrm{train\&val}}\setminus\left(\mathcal{\cup}_{i=6-|\rho|)}^{5}\mathcal{E}_{i}\right)&\mathrm{if~\rho<0}\end{cases}$ (11)
Where $\cup$ indicates set union and $\setminus$ indicates set difference. $\rho\in\{-4,-3,-2,-1,0,1,2,3,4,5\}$ reflects level of B-I variation. $\rho>0$ indicates adding set $\Delta^{+}$ to $Z_{\mathrm{test}}$, $\rho<0$ indicates removing set $\Delta^{-}$ from $Z_{\mathrm{test}}$ and $\rho=0$ indicates $Z_{\mathrm{test}}$ invariability. The well-trained bundle models predict user-bundle interactions $X$ on $Z_{\mathrm{test}}$. The processing of the B-I affiliation is shown in Figure 3. In this way, we can control the degree of B-I affiliatons variation and simulate the scenario of B-I dynamic variability. Recall@K and NDCG@K are used as the evaluation metrics, where K $\in\{10,20,40,80\}$.
$\bullet$ Hyperparameter Settings The hyperparameters of six backbone models are set by following their publications and RDiffBR is optimized with the Adam optimizer with a learning rate of 0.001. The time steps $T$ and sampling steps $T^{\prime}$ are set to values ranging from [1,200]. The loss balance factor $\lambda$ is chosen from (0,5), while the residual connection weight $\delta$ is selected from (0,1). The number of MLP layers is selected within 1, 2, 4, 8. Its hidden sizes are from [8,4096]. Details of other hyperparameters can be found in our released code. All the models are trained using two GTX 4090.
[…]
Overall Performance (RQ1)
The following observations are from Table 2.
@@
$\bullet$ The performance of BR backbones when meeting item-level dynamic variability. When $\rho<0$ ($\mathbf{Z_{test}}$ removes $\Delta^{-}$), the performance of bundle models is generally lower than that under $\rho=0$ in most cases. The greater the reduction in B-I, the more significant the performance degradation. When $\rho>0$ ($\mathbf{Z_{test}}$ adds $\Delta^{+}$), some models exhibit performance decline, others maintain comparable performance, while a few show improvement. This indicates that BR models possess capability to handle B-I variations to some extent, yet they tend to fail under drastic reductions of B-I. Notably, HyperMBR performs the worst, where both B-I adding and removing frequently lead to performance deterioration. This phenomenon may stem from hyperbolic space embeddings being more sensitive to distance errors compared to Euclidean space. In contrast, models like MIDGN, CrossCBR, and MultiCBR exhibit relatively stable performance, potentially attributed to their adoption of contrastive learning that leverages multi-view information through cooperative learning. Additionally, BundleGT demonstrates enhanced adaptability for dynamic B-I variations by modeling B-I affiliation via attention mechanisms.
@@
Ablation Study (RQ2)
@@
[…]
@@
Table 3: Ablation study where bold indicates the best method, and underlined indicates the second best method($\rho=-3,3$).
Figure 5: Blue is distribution $E^{IL}_{B}$ under $\rho=0$. Red is distribution $E^{IL}_{B}$ under $\rho=-3$.
Figure 6: Blue is distribution $E^{IL}_{B}$ under $\rho=0$. Red is distribution $E^{IL}_{B}$ under $\rho=3$.
@@
From this table, we can draw the following conclusions: In most cases, the performance of using RDiffBR is superior to that of not using it or using RDiffBR w/o R. Using RDiffBR w/o R may lead to worse performance than RDiffBR, and in some cases it may even be worse than only using backbone BR model. Figure 6 shows the visualization results of CrossCBR, CrossCBR+RDiffBR and CrossCBR+RDiffBR w/o R under B-I varation levels $\rho=3$ for item-level bundle embeddings $E^{IL}_{B}$ in Youshu dataset. It can be observed that $E^{IL}_{B}$ generated by CrossCBR after B-I change is more dispersed than original $E^{IL}_{B}$ and more compact $E^{IL}_{B}$ distribution which retains the basic shape can be obtained after using CrossCBR+RDiffBR. This indicates that $E^{IL}_{B}$ of the BR model cannot effectively represent the bundle theme distribution due to the B-I change and RDiffBR enables the model to obtain more bundle information at the item-level rather than overfitting to the existing b-I affiliations. While using CrossCBR+RDiffBR w/o R, $E^{IL}_{B}$ distribution changes little so the effect is limited, which is consistent with the experimental results obtained in Table 3. Overall, RDiffBR outperforms backbone models, highlighting the effectiveness of RDiffBR. Residual connection plays an important role in RDiffBR, and merely using a diffusion model cannot generate better item-level bundle embeddings (The performance of $\rho=-3$ is consistent with $\rho=3$, see details in extended version).
@@
Although different B-I affiliations are used during model testing, the bundle theme remains, which means items within a bundle remain to follow the similar distribution before and after the change. This can also be verified by Figure 6 that the shapes of two $E^{IL}_{B}$ distributions under different B-I variation $\rho$ are similar. Therefore, with using RDiffBR, our task can be approached without considering the Out-Of-Distribution (OOD) where the distribution of the training data and the test data is inconsistent. In extreme cases where the distribution of $E^{IL}_{B}$ changes completely, the problem becomes equivalent to cold-start bundle recommendation (cold_start_1; cold_start_2) that the original bundle id cannot be directly associated.
Impact of Hyperparameters
Figure 7: The impact of loss balance factor $\lambda$ .
Figure 8: The impact of residual weight $\delta$ .
@@
In order to achieve a more detailed understanding, two key hyperparameters are studied, i.e., the loss balance factor $\lambda$ and the residual weight $\delta$ . The results of our hyperparameter study are presented in Figure 8-8.
@@
$\bullet$ Loss balance factor $\lambda$. The loss balance factor $r$ affects the performance of RDiffBR, which adjusts the emphasis between the recommendation task and the reconstruction task. A high value of $r$ may neglect the core recommendation and thereby compromise the overall performance.
@@
$\bullet$ Residual weight $\delta$. The appropriate selection of residual weights is crucial for embedding generation, this is because too high or too low values will lead to a performance decline.
Case Study
@@
Figure 9: A case study for Youshu under $\rho=-3$. The highlighted items are those before the bundle change.
@@
Figure 10: A case study for Youshu under $\rho=3$. The highlighted items are those before the bundle change.
@@
In this section, we analyze the real data instances as show in Figure 9. We randomly select a bundle from Youshu and calculate cosine similarity between item-level bundle embeddings $E^{IL}_{B}$ and all items under B-I varation levels $\rho=-3$. We listed the top 5 items with the highest similarity scores and their corresponding similarity values. It can be seen that the similarity of items (marked in red) that appeared within the bundle will be higher when only using CrossCBR, resulting in insufficient generalization ability. When using RDiffBR w/o R, the model cannot well capture the main items of the bundle. When using RDiffBR, the model can not only express the main items of the bundle but also introduce other items based on the theme to enhance the generalization ability.
Training Efficiency Study (RQ3)
@@
In order to evaluate the additional computing time of RDiffBR, we compared the one-epoch training time (seconds) of using BR model and BR model+RDiffBR on four datasets (The training rounds are the same for both modes). We recorded five consecutive training epochs and averaged them to obtain the one-epoch training time, as shown in Table 4. It shows that RDiffBR only increased training time about 4%.
[…]
BGCN 1.46 23.5 68.3 1.07
+RDiffBR 1.51 24.5 71.4 1.09
Extra cost 3.42% 4.25% 4.53% 1.86%
CrossCBR 1.06 8.70 45.4 1.10
+RDiffBR 1.11 9.10 46.2 1.14
Extra cost 4.71% 4.59% 1.76% 3.63%
BundleGT 22.7 426.2 118.4 1.23
+RDiffBR 23.3 434.8 121.1 1.29
Extra cost 2.64% 2.01% 2.28% 4.87%